A village primary school has 120 pupils. 30 of them walk to school, 48 are brought by car, 36 come on the school bus and 6 of them cycle. Convert these figures to percentages of the total.
Please help me.

Answer by Alan3354(69443)

A village primary school has 120 pupils. 30 of them walk to school, 48 are brought by car, 36 come on the school bus and 6 of them cycle. Convert these figures to percentages of the total.
----------
30/120 = 1/4
(1/4)*100% = 25% walk
------
Do the others the same way.